OpenStudy (anonymous):

what squared=(cscx-1)(cscx+1)

OpenStudy (anonymous):

u mean we have to find x for which \[x^2=(cscx-1)(cscx+1)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x^2=\cot^2x\]

OpenStudy (anonymous):